If you spin the spinner 100 times, how many times do you expect it to stop on 8? The spinner is divided into 8 parts.
Answer by Boreal(15235)

E(X)=X*p(x)=100*0.125=12.5 times. Fractions are accepted as expected values. Realistically, would one expect 12 or 13 times.